# Efficient and Accurate Optimal Transport with Mirror Descent and Conjugate Gradients

### v3

@@ 1 INTRODUCTION
    We show that MDOT can compute high precision, feasible solutions and its performance can be boosted by adopting a specialized GPU-parallel conjugate gradients (CG) algorithm developed here (Alg. 2); this method is highly competitive in practice, as we show empirically (Figs. 4, 5, 8-17).
@@ 2.1 Optimal Transport
Given a cost matrix ${C\in{[0,1]}^{n\times n}}$, where $C_{ij}$ is the transportation cost between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ particles, we study the EOT problem given by: {mini}[2] P ∈U(r, c)⟨P, C ⟩- 1γH(P), where $\gamma>0$. Here, the regularization weight $\gamma^{-1}$ is called temperature. The Lagrangian of (2.1) is strictly convex in $P$, which renders the solution $P^{*}(\gamma)$ unique. $P^{*}(\gamma)$ converges to a solution of the unregularized OT problem as $\gamma\rightarrow\infty$ and admits the following form (Cuturi, 2013):
$P_{ij}({\bm{u}},{\bm{v}};\gamma)=\exp\{u_{i}+v_{j}-\gamma C_{ij}\},$ (1)
where ${{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}$. An optimal pair $({\bm{u}},{\bm{v}})$ minimizes the following convex dual problem (Lin et al., 2019): {mini}[2] u, vR^ng(u, v; γ) = 1^⊤P(u, v; γ) 1 - ⟨u, r⟩- ⟨v, c⟩, where $\nabla_{\bm{u}}g={\bm{r}}(P)-{\bm{r}}$ and $\nabla_{\bm{v}}g={\bm{c}}(P)-{\bm{c}}$. Given some initial $({\bm{u}},{\bm{v}})$, solving (2.1) amounts to a KL projection of $P({\bm{u}},{\bm{v}};\gamma)$ onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$. That is, $P({\bm{u}}^{*},{\bm{v}}^{*};\gamma)=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\big{(}P|P({\bm{u}},{\bm{v}};\gamma)\big{)}$; see derivation in Appx. A.1.
The Sinkhorn-Knopp (SK) algorithm (see Alg. 4 in Appx. A) can be used to solve (2.1) and carry out an approximate projection of $P({\bm{u}},{\bm{v}};\gamma)$ onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$. Sinkhorn updates guarantee convergence to dual-optimal variables as the number of iterations $k\rightarrow\infty$ (Sinkhorn & Knopp, 1967; Sinkhorn, 1967; Franklin & Lorenz, 1989; Knight, 2008). Dvurechensky et al. (2018) showed that SK can be used to compute a solution $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ satisfying $\langle P-P^{*},C\rangle\leq\varepsilon$ with complexity $\widetilde{O}(n^{2}/\varepsilon^{2})$, where $P^{*}$ is an optimal solution of the unregularized OT problem. In particular, one first minimizes the dual objective until the $L_{1}$-norm of its gradient is below a prescribed threshold, then applies the rounding algorithm of Altschuler et al. (2017) on the infeasible plan given by (1) to obtain $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ with an upper bound on the primal cost increase.
@@ 2.2 Mirror Descent
Rather than the entropic regularized problem (2.1), we will approximately solve a sequence of relative-entropy regularized problems as in mirror descent, where each problem has the form (Nemirovski & Yudin, 1983):
$P^{(t+1)}=\operatorname*{arg\,min}_{P\in{\mathcal{F}}\cap{\mathcal{D}}}\{\langle\nabla_{P}f(P^{(t)}),P\rangle+\frac{1}{\Delta^{(t)}}D_{h}(P|P^{(t)})\}.$ (2)
Here $f:{\mathcal{D}}\rightarrow\mathbb{R}$ is the objective, ${\mathcal{F}}$ is a given feasible set, $\Delta^{(t)}{>}0$ is the step size and the Bregman divergence
$D_{h}(Q|P)=h(Q)-h(P)-\langle\nabla h(P),Q-P\rangle$ (3)
given a strictly convex and differentiable function ${h:{\mathcal{D}}\rightarrow\mathbb{R}}$ called the mirror map. Equivalent to (2) is:
$\hat{P}^{(t+1)}$ $=\nabla h^{-1}\left(\nabla h(P^{(t)})-\Delta^{(t)}\nabla f(P^{(t)})\right)$ (4)
$P^{(t+1)}$ $=\operatorname*{arg\,min}_{P\in{\mathcal{F}}\cap{\mathcal{D}}}D_{h}(P|\hat{P}^{(t+1)}).$ (5)
Here, (4) takes a gradient step in the dual space and maps the new point back onto the primal space via $\nabla h^{-1}$, while (5) defines a Bregman projection of $\hat{P}^{(t+1)}$ onto the feasible set ${\mathcal{F}}$ in the primal space.
Throughout, we only use the negative entropy mirror map ${h(P)=-H(P)}$ in domain ${{\mathcal{D}}=\mathbb{R}^{n\times n}_{\geq 0}}$, which yields ${D_{h}(Q|P)=D_{\mathrm{KL}}(Q|P)}$. Di Marino & Gerolin (2020) provide a treatment of more general divergence-regularized optimal transport problems (using mirror maps besides negative Shannon entropy), albeit running only a single step of (2) with $P^{(0)}\in\mathbb{R}^{n\times n}_{\geq 0}$ set to the independence coupling ${\bm{r}}{\bm{c}}^{\top}$, ${\mathcal{F}}={\mathcal{U}}({\bm{r}},{\bm{c}})$ and objective $f(P)=\langle P,C\rangle$; we consider multiple steps.
@@ 3 Related Work
Application of mirror descent to the primal of the OT problem has also been considered. Yang & Toh (2022) discuss an algorithm similar to MDOT, although their approach differs from ours in a number of ways. They require a rounding procedure onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$ after each mirror descent step, and verify complicated stopping criteria for all Bregman projections; we did not find these to be necessary experimentally or mathematically. Recently, Ballu & Berthet (2023) introduced Mirror Sinkhorn (MSK), which also takes gradient steps in the dual space as in (4), but instead of approximately projecting onto the feasible set ${\mathcal{U}}({\bm{r}},{\bm{c}})$ as in (5) (as we do here), they alternately project onto ${\mathcal{U}}(\cdot,{\bm{c}})$ and ${\mathcal{U}}({\bm{r}},\cdot)$ via Sinkhorn updates, satisfying only half of the equality constraints at a time. Our experiments in Sec. 5 suggest that this approach is efficient only in the low precision regime. Furthermore, MSK requires maintaining a running average of the transport plan at each iteration, precluding a straightforward $O(n)$ memory implementation. One could, in principle, store all past dual variables and recompute each plan to form the average at each step, but this would require $O(nT)$ memory and $O(n^{2}T)$ recomputation time at iteration $T$, which is typically impractical. In contrast, all algorithms presented here admit $O(n)$ memory implementations, assuming individual cost matrix entries can be computed on-the-fly in $O(1)$ time. Xie et al. (2020) previously proposed an algorithm (IPOT) similar to MSK with a fixed, even number of Sinkhorn updates (usually 2) following temperature updates; we omitted additional empirical comparison to IPOT given its similarity with MSK. Alg. 3.5 of Feydy (2020) is also similar to these algorithms in spirit and is discussed thoroughly in Sec. 5.2. As discussed in detail in Sec. 4.1, well-known $\varepsilon$-scaling strategies are also closely related (Kosowsky & Yuille, 1994; Schmitzer, 2019).
An alternative line of acceleration research focuses on multi-scale strategies, which employ clustering or grid-based methods to solve a series of coarse-to-fine OT problems and are sometimes combined with $\varepsilon$-scaling (Schmitzer, 2016, 2019; Feydy, 2020). These are known to provide performance gains when the marginals are defined over well-clustered particles or in low-dimensional event spaces (Peyré et al., 2019). Lastly, in a similar spirit to our use of non-linear CG here, curvature-aware convex optimization techniques such as L-BFGS have also been considered for OT, e.g., Mérigot (2011); Blondel et al. (2018); however, scalability, precision and better performance than SK on GPUs has not been demonstrated simultaneously to our knowledge. Tang et al. (2024) recently adopted Newton’s method with Hessian sparsification to efficiently use second order information, but their key sparsification strategy is maximally utilized only on CPUs.
@@ 4.1 Temperature Annealing as Mirror Descent
The OT objective $\langle P,C\rangle$ has a constant gradient ${\nabla_{P}\langle P,C\rangle=C}$. Given step sizes $\Delta_{\gamma}^{(t)}>0$ at time $t\geq 0$, mirror descent iterates with the negative entropy mirror map $h(P)=-H(P)$ are thus given by (cf. 2):
$P^{(t+1)}=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}\left\{\langle P,C\rangle+\frac{1}{\Delta_{\gamma}^{(t)}}D_{\mathrm{KL}}(P|P^{(t)})\right\}.$ (6)
A useful convention here is to take $P^{(0)}={\bm{r}}{\bm{c}}^{\top}$, the maximum entropy transport plan in ${\mathcal{U}}({\bm{r}},{\bm{c}})$. However, as the second item of the next proposition, which describes the properties of $P^{(t)}$ computed via (6), suggests, this initialization is not strictly necessary and any rank-1 matrix with positive entries may be selected.
@@
Suppose $P^{(0)}\in\mathbb{R}^{n\times n}_{>0}$ is rank-1 and $P^{(t)}$ are computed via (6) for $t\geq 1$. Let ${\gamma^{(0)}=0}$ and ${\gamma^{(t+1)}=\gamma^{(t)}}+\Delta_{\gamma}^{(t)}$, which together imply $\gamma^{(t+1)}=\sum_{t^{\prime}=0}^{t}\Delta_{\gamma}^{(t^{\prime})}$. Given ${P^{*}\in{\mathcal{U}}({\bm{r}},{\bm{c}})}$, a minimizer of $\langle P,C\rangle$, and $H_{\min}({\bm{r}},{\bm{c}})\coloneqq\min\big{(}H({\bm{r}}),H({\bm{c}})\big{)}$, the following are true:
@@
    $P^{(t)}=P^{*}(\gamma^{(t)})$, where $P^{*}(\gamma^{(t)})$ is the unique solution of (2.1) for $\gamma=\gamma^{(t)}$.
@@
    Let ${\widetilde{P}^{(t)}\in\mathbb{R}^{n\times n}_{>0}}$ be any matrix of the form $\widetilde{P}_{ij}^{(t)}=\exp\{u_{i}+v_{j}-\gamma^{(t)}C_{ij}\}$ for some ${\bm{u}},{\bm{v}}\in\mathbb{R}^{n}$. In addition to (6), we also have $P^{(t+1)}=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}\left\{\langle P,C\rangle+\frac{1}{\Delta_{\gamma}^{(t)}}D_{\mathrm{KL}}(P|\widetilde{P}^{(t)})\right\}$.
@@
    $\langle P^{(t)}-P^{*},C\rangle\leq H_{\min}({\bm{r}},{\bm{c}})\big{/}\gamma^{(t)}$.
@@
The equivalence of $P^{(t)}$ to the solution of (2.1) can be seen as follows. By setting the partial derivatives (with respect to $P$) of the Lagrangian of (6) to zero, we find that $P^{(t+1)}$ must have the form
$P_{ij}^{(t+1)}=P^{(t)}_{ij}\exp\{u_{i}+v_{j}-\Delta_{\gamma}^{(t)}C_{ij}\}.$ (7)
These relations allow us to unroll the recursion to get $P^{(t+1)}$ as follows (using the fact that $P^{(0)}$ is rank-1):
$P_{ij}^{(t+1)}=\exp\{u^{*}_{i}(\gamma^{(t)})+u_{i}+v^{*}_{j}(\gamma^{(t)})+v_{j}-\gamma^{(t+1)}C_{ij}\}.$ (8)
Thus, $P$ in (7-8) has the same form as (1). That is, the mirror descent procedure in (6) mapped to the dual space can be understood as a temperature annealing strategy (called $\varepsilon$-scaling by Schmitzer (2019)), in which dual-optimal ${\bm{u}},{\bm{v}}$ are found at each value of the decaying temperature $1/\gamma^{(t)}$. Notice here that due to strong duality, optimal ${\bm{u}}^{*}(\gamma^{(t)}),{\bm{v}}^{*}(\gamma^{(t)})$ are minimizers of the EOT dual objective $g({\bm{u}},{\bm{v}};\gamma^{(t)})$ in (2.1). The sequence of dual problems correspond to a sequence of Bregman projections like (5) in the primal space.
The second item in Prop. 4.1 shows that in fact, intermediate dual problems need not be solved exactly to optimality for this procedure to ultimately arrive at $P^{*}(\gamma)$. In other words, we can use approximations for ${\bm{u}}^{*}(\gamma^{(t)})$ and ${\bm{v}}^{*}(\gamma^{(t)})$ in (8) and solve for the update vectors ${\bm{u}}$ and ${\bm{v}}$ to approximate $P^{(t+1)}$. Notice that an initialization $({\bm{u}}^{(0)},{\bm{v}}^{(0)})$ here corresponds to $\widetilde{P}^{(0)}_{ij}=\exp\{u^{(0)}_{i}+v^{(0)}_{j}\}$ with $\gamma^{(0)}=0$, which is why any rank-1 $P^{(0)}\in\mathbb{R}^{n\times n}_{>0}$ suffices. These observations will form the basis of MDOT (Alg. 1).
The third item in Prop. 4.1 bounds the excess cost of the primal objective at a given step $t\geq 1$ in terms of the entropies of the marginals ${\bm{r}},{\bm{c}}$. It should be compared to the more standard upper bound $\gamma^{-1}\log n$ used in prior work (Altschuler et al., 2017; Dvurechensky et al., 2018; Lin et al., 2019). [^2] This $\log n$ term appears in the time complexity of various algorithms, but is often hidden in $\widetilde{O}$-notation. Note $H_{\min}({\bm{r}},{\bm{c}})\leq\log n$ for ${\bm{r}},{\bm{c}}\in\Delta_{n}$ and therefore this is a tighter bound. The final item in Prop. 4.1 shows the one-step improvement in the linear objective with equality; this is in contrast to the more standard analysis of mirror descent procedures where the improvement is bounded above with an inequality. See Appx. A.2 for proofs.

### v4

@@ 1 INTRODUCTION
    We show that MDOT can compute high precision, feasible solutions and its performance can be boosted by adopting a specialized GPU-parallel conjugate gradients (CG) algorithm developed here (Alg. 2); this method is highly competitive in practice, as we show empirically (Figs. 5, 6, 10-19).
@@ 2.1 Optimal Transport
Given a cost matrix ${C\in{[0,1]}^{n\times n}}$, where $C_{ij}$ is the transportation cost between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ particles, we study the EOT problem: {mini}[2] P ∈U(r, c)⟨P, C ⟩- 1γH(P), where $\gamma>0$. Here, the regularization weight $\gamma^{-1}$ is called temperature. The Lagrangian of (2.1) is strictly convex in $P$, which renders the solution $P^{*}(\gamma)$ unique. $P^{*}(\gamma)$ converges to a solution of the unregularized OT problem as $\gamma\rightarrow\infty$ and admits the following form (Cuturi, 2013):
$P_{ij}({\bm{u}},{\bm{v}};\gamma)=\exp\{u_{i}+v_{j}-\gamma C_{ij}\},\vspace{-5pt}$ (1)
where ${{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}$. An optimal pair $({\bm{u}},{\bm{v}})$ minimizes the following convex dual problem (Lin et al., 2019): {mini}[2] u, vR^ng(u, v;  γ, r, c) = ∑_ijP_ij(u, v; γ) - ⟨u, r⟩- ⟨v, c⟩, where $\nabla_{\bm{u}}g={\bm{r}}(P)-{\bm{r}}$ and $\nabla_{\bm{v}}g={\bm{c}}(P)-{\bm{c}}$. The gradient norm naturally measures the constraint violation; $L_{1}$ norm is typically used to monitor convergence as it relates to the total variation metric between probability distributions (Altschuler et al., 2017). The Sinkhorn-Knopp (SK) algorithm (see Alg. 4 in Appx. A) can be used to minimize $g$ with guaranteed convergence to dual-optimal variables as the number of iterations $k\rightarrow\infty$ (Sinkhorn & Knopp, 1967; Sinkhorn, 1967; Franklin & Lorenz, 1989; Knight, 2008). Dvurechensky et al. (2018) showed that SK can be used to compute a solution $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ satisfying $\langle P-P^{*},C\rangle\leq\varepsilon$ with complexity $\widetilde{O}(n^{2}/\varepsilon^{2})$, where $P^{*}$ is an optimal solution of the unregularized OT problem. In particular, one first minimizes the dual objective until the $L_{1}$ norm of its gradient is below a prescribed threshold, then applies the rounding algorithm of Altschuler et al. (2017) on the infeasible plan given by (1) to obtain $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ with an upper bound on the primal cost increase.
@@ 2.2 Mirror Descent
Consider a constrained convex minimization problem $\min_{P\in{\mathcal{F}}}f(P)$, where the convex objective ${f:{\mathcal{D}}\rightarrow\mathbb{R}}$ is defined over some domain ${\mathcal{D}}$, and the feasible set ${{\mathcal{F}}\subseteq{\mathcal{D}}}$, e.g., in OT, we take ${f(P)=\langle P,C\rangle}$, ${{\mathcal{F}}={\mathcal{U}}({\bm{r}},{\bm{c}})}$ and ${{\mathcal{D}}=\mathbb{R}^{n\times n}_{\geq 0}}$. Mirror descent (MD) method of Nemirovski & Yudin (1983) generalizes projected gradient descent (GD) by first choosing a strictly convex function $h:{\mathcal{D}}\rightarrow\mathbb{R}$ called the mirror map, which induces a Bregman divergence ${D_{h}(P|Q)\geq 0}$ between points $P,Q\in{\mathcal{D}}$ as the difference between $h(P)$ and its first order approximation around $Q$:
$D_{h}(P|Q)=h(P)-h(Q)-\langle\nabla h(Q),P-Q\rangle.$ (2)
Then, the variable $P^{(t)}$ is updated as follows (Bubeck, 2015):
where ${\Delta^{(t)}>0}$ is the step size. For $h(P)=\left\lVert P\right\rVert_{2}^{2}/2$, projected GD is recovered as a special case since ${\nabla h(P)=P}$ and $D_{h}(P|Q)=\left\lVert P-Q\right\rVert^{2}_{2}/2$. For certain problem geometries (e.g., if ${\mathcal{D}}$ is the simplex), choosing a different $h$ provably accelerates convergence. The idea is that in (3a), ${\nabla h:{\mathcal{D}}\rightarrow{\mathcal{D}}^{*}}$ maps primal variables $P$ to a dual space ${\mathcal{D}}^{*}$, where applying a gradient update better aligns with the local curvature of the underlying geometry (Bubeck, 2015). Once the updated point is mapped back to primal space ${\mathcal{D}}$ via $\nabla h^{-1}:{\mathcal{D}}^{*}\rightarrow{\mathcal{D}}$, (3b) performs a Bregman projection onto the feasible set ${\mathcal{F}}$. Plugging (3a) into (3b) and rearranging yields an equivalent form with another interpretation (Beck & Teboulle, 2003):
$P^{(t+1)}=\operatorname*{arg\,min}_{P\in{\mathcal{F}}\cap{\mathcal{D}}}\{\langle\nabla_{P}f(P^{(t)}),P\rangle+\frac{1}{\Delta^{(t)}}D_{h}(P|P^{(t)})\}.\vspace{-15pt}$ (4)
Each MD step (4) seeks an update jointly maximizing the inner product with the steepest descent direction, $-\nabla_{P}f(P^{(t)})$, subject to (i) a penalty for diverging from the current point $P^{(t)}$ and (ii) feasibility constraints.
A common choice is the negative entropy mirror map ${h(P)=\langle P,\log P\rangle}$ in domain ${{\mathcal{D}}=\mathbb{R}^{n\times n}_{\geq 0}}$, which yields ${D_{h}(P|Q)=D_{\mathrm{KL}}(P|Q)}$. Throughout this work, we only use this mirror map. [^1] See Di Marino & Gerolin (2020) for other divergence-regularized optimal transport problems (using mirror maps besides negative Shannon entropy). They consider a single step of (4), while we focus on multiple steps. Hence, we write (3) explicitly for this case, where ${\nabla h(P)=\bm{1}+\log P}$ and $\nabla h^{-1}(R)=\exp\{R-\bm{1}\}$,
Here, we re-weight $P^{(t)}$ entrywise by a factor of $\exp\{-\Delta^{(t)}\nabla f(P^{(t)})\}$, followed by a KL projection onto ${\mathcal{F}}$, e.g., if ${\mathcal{F}}$ is the probability simplex, (5b) is a simple renormalization $P\mapsto P/\left\lVert P\right\rVert_{1}$. We leverage the properties of this mirror map for the case ${\mathcal{F}}={\mathcal{U}}({\bm{r}},{\bm{c}})$ to design MDOT in Sec. 4.
@@ 3 Related Work
Application of mirror descent to the primal of the OT problem has also been considered. Yang & Toh (2022) propose a more general inexact mirror descent algorithm (iBPPA), which they apply to the OT problem. Our approach differs from theirs in several ways, most notably in how we decide to terminate the Bregman projection inner loop to solve (5b) inexactly (details in Section 4.2). Recently, Ballu & Berthet (2023) introduced Mirror Sinkhorn (MSK), which also takes gradient steps in the dual space as in (5a), but instead of approximately projecting onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$ as in (5b) (as we do here), they alternately project onto ${\mathcal{U}}(\cdot,{\bm{c}})$ and ${\mathcal{U}}({\bm{r}},\cdot)$ via Sinkhorn updates, satisfying only half of the marginal constraints at a time. Our experiments in Sec. 5 suggest that this approach is efficient only in the low precision regime. Furthermore, MSK requires maintaining a running average of the transport plan at each iteration, precluding a straightforward $O(n)$ memory implementation. In contrast, all algorithms presented here admit $O(n)$ memory implementations, assuming individual cost matrix entries can be computed on-the-fly in $O(1)$ time. Xie et al. (2020) previously proposed an algorithm (IPOT) similar to MSK with a fixed, even number of Sinkhorn updates (usually 2) following temperature updates. Alg. 3.5 of Feydy (2020) is also similar to these algorithms in spirit and is discussed thoroughly in Sec. 5.2. As discussed in detail in Sec. 4.1, well-known $\varepsilon$-scaling strategies are also closely related (Kosowsky & Yuille, 1994; Schmitzer, 2019). Similar ideas have also been applied to the Wasserstein barycenter problem (Gramfort et al., 2015; Xie et al., 2020), which is left outside the scope of this work.
An alternative line of acceleration research focuses on multi-scale strategies, which employ clustering or grid-based methods to solve a series of coarse-to-fine OT problems and are sometimes combined with $\varepsilon$-scaling (Schmitzer, 2016; 2019; Feydy, 2020). These are known to provide performance gains when the marginals are defined over well-clustered particles or in low-dimensional event spaces (Peyré et al., 2019). Lastly, in a similar spirit to our use of non-linear CG here, curvature-aware convex optimization techniques such as L-BFGS have also been considered for OT, e.g., Mérigot (2011); Blondel et al. (2018); however, scalability, precision and better performance than SK on GPUs has not been demonstrated simultaneously to our knowledge. Tang et al. (2024) recently adopted Newton’s method with Hessian sparsification to efficiently use second order information, but their key sparsification strategy is maximally utilized only on CPUs.
@@ 4.1 Temperature Annealing as Mirror Descent
The OT objective $\langle P,C\rangle$ has a constant gradient ${\nabla_{P}\langle P,C\rangle=C}$. Given step sizes $\Delta^{(t)}>0$ at time $t\geq 0$, we obtain mirror descent iterates (using the negative entropy mirror map) from (5)
$P^{(t+1)}=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P^{(t)}\odot\exp\{-\Delta^{(t)}C\}\Big{)}.\vspace{-15pt}$ (6)
Comparing the conditioning term $P^{(t)}\odot\exp\{-\Delta^{(t)}C\}$ above with the Lagrangian closed form solution to the EOT problem in (1), namely $P({\bm{u}},{\bm{v}};\gamma)=P({\bm{u}}\bm{1}^{\top}+\bm{1}{\bm{v}}^{\top}-\gamma C)$, suggests considering an MD step starting with a $P^{(t)}$ of this same form. Specifically, we find the following result (proved in Appx. A.1).
Lemma 4.1.
Given any initial ${{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}$, $\gamma\geq 0$ and $\Delta_{\gamma}>0$, we have
$P({\bm{u}}^{*},{\bm{v}}^{*};\gamma+\Delta_{\gamma})=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P({\bm{u}},{\bm{v}};\gamma)\odot\exp\{-\Delta_{\gamma}C\}\Big{)},$ (7)
where ${\bm{u}}^{*},{\bm{v}}^{*}\in\operatorname*{arg\,min}g(\gamma+\Delta_{\gamma},{\bm{r}},{\bm{c}})$.
That is, if we begin an MD step for the OT problem with any $P^{(t)}=P({\bm{u}},{\bm{v}};\gamma^{(t)})$ (i.e., in the Lagrangian form but not necessarily optimal for the EOT problem), then the next MD iterate is the unique solution to the EOT dual problem at $\gamma^{(t+1)}=\gamma^{(t)}+\Delta_{\gamma}^{(t)}$, namely $P^{(t+1)}=P^{*}(\gamma^{(t+1)})$. Therefore, by induction, if we begin the MD iteration (6) for the OT problem with such an initial $P^{(0)}$, then the MD iterates $P^{(t)}$ trace solutions of the EOT problems at increasing values of $\gamma^{(t)}$.
For initialization, recall that the EOT solution at $\gamma{=}0$ is the maximum entropy coupling, ${\bm{r}}{\bm{c}}^{\top}=\lim_{\gamma\rightarrow 0^{+}}\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}\langle P,C\rangle-\frac{1}{\gamma}H(P)$. Thus, we can initialize $\gamma^{(0)}=0$ and $P^{(0)}=P(\log{\bm{r}},\log{\bm{c}};0)={\bm{r}}{\bm{c}}^{\top}$. Alternatively, in view of Lemma 4.1, it would suffice to use $P^{(0)}=P({\bm{u}},{\bm{v}};0)$ for any ${\bm{u}},{\bm{v}}\in\mathbb{R}^{n}$ at $\gamma^{(0)}=0$. That is, $P^{(0)}$ can be any rank-1 matrix in >$\mathbb{R}^{n\times n}_{>0}$.
Finally, note that for the purpose of computing $P^{(t+1)}$ in (6), Lemma 4.1 ensures that any values ${\bm{u}}^{(t)}$ and ${\bm{v}}^{(t)}$ can serve to provide a suitable $P^{(t)}=P({\bm{u}}^{(t)},{\bm{v}}^{(t)};\gamma^{(t)})$. Therefore, when implementing the MD iteration in (6) we do not need to converge to exact optimality each step (see Fig. 1). We formalize this discussion and provide further insight in the following proposition.
@@
Let ${\gamma^{(0)}=0}$ and ${\gamma^{(t+1)}=\gamma^{(t)}}+\Delta_{\gamma}^{(t)}$, which together imply $\gamma^{(t+1)}=\sum_{t^{\prime}=0}^{t}\Delta_{\gamma}^{(t^{\prime})}$. Suppose $P^{(0)}\in\mathbb{R}^{n\times n}_{>0}$ is rank-1 and $P^{(t)}$ are computed via (6) for $t\geq 0$. The following are true.
@@
    $P^{(t+1)}=P^{*}(\gamma^{(t+1)})$, i.e., the solution of the EOT problem (2.1) at $\gamma^{(t+1)}$.
@@
    Given any ${\bm{u}},{\bm{v}}\in\mathbb{R}^{n}$, $P^{(t+1)}=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}(P|P({\bm{u}},{\bm{v}};\gamma^{(t+1)}))$.
@@
    $\langle P^{(t)}-P^{*},C\rangle\leq H_{\min}({\bm{r}},{\bm{c}})\big{/}\gamma^{(t)}$, where ${P^{*}\in\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}\langle P,C\rangle}$ and $H_{\min}({\bm{r}},{\bm{c}})\coloneqq\min\big{(}H({\bm{r}}),H({\bm{c}})\big{)}$.
@@
A detailed proof is deferred to Appx. A.2. Next we comment on each statement in the order presented.
1. proves the equivalence of MD iterates to a sequence of solutions to EOT problems with decreasing temperature, and therefore connects MD to temperature annealing.
2. shows that MD iterates $P^{(t+1)}$ derived from a rank-1 initialization can be written independently of the previous solution $P^{(t)}$ (cf. 6), but rather only as a function of $\gamma^{(t+1)}$. This means that we need not start from an exact minimizer of the previous KL projection problem to reach the correct solution of the new problem. Given this result and Lemma 4.1, MD reduces in the dual space to a numerical continuation method (Allgower & Georg, 2003), in which we numerically trace the curve ${\bm{u}}^{*}(\gamma),{\bm{v}}^{*}(\gamma)$: [^2] Slight abuse of terminology: in fact, there is a space of optimal curves, since $g({\bm{u}},{\bm{v}};\gamma)=g({\bm{u}}{+}\delta\bm{1},{\bm{v}}{-}\delta\bm{1};\gamma)$ for any finite $\delta$.
${\bm{u}}^{(t)},{\bm{v}}^{(t)}\approx\operatorname*{arg\,min}_{{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}g({\bm{u}},{\bm{v}};\gamma^{(t)},{\bm{r}},{\bm{c}}),$ (8)
where each problem is initialized with some guess ${\bm{u}}^{\prime},{\bm{v}}^{\prime}\in\mathbb{R}^{n}$. This is the key idea behind MDOT.
3. bounds the primal optimality gap at a given step $t\geq 1$ in terms of the entropies of the marginals ${\bm{r}},{\bm{c}}$. Since $\min(H({\bm{r}}),H({\bm{c}}))\leq\log n$ for ${\bm{r}},{\bm{c}}$ on the ${(n{-}1)}$-simplex, this is a tighter bound than the standard upper bound $\gamma^{-1}\log n$ used in prior work (Altschuler et al., 2017; Dvurechensky et al., 2018; Lin et al., 2019). [^3] This $\log n$ term appears in the time complexity of various algorithms, but is hidden in $\widetilde{O}$-notation.
4. shows the one-step reduction in transport cost with equality; this is in contrast to the more standard analysis of MD where the improvement is bounded with an inequality.